### v1

@@ Abstract
Keywords:
Agentic UAVs Large Language Models Tool-Calling Cognitive UAV Autonomy, Autonomous Aerial Systems
@@ 1 Introduction
Unmanned Aerial Vehicles (UAVs) are rapidly transitioning from remotely piloted platforms to autonomous agents capable of performing increasingly complex tasks in logistics, agriculture, defense, and disaster response. This transformation parallels advances in artificial intelligence (AI), where systems have evolved from rule-based automation to data-driven and adaptive reasoning. Despite progress in perception, planning, and control, UAV autonomy remains predominantly narrow—optimized for specific tasks but lacking the general-purpose intelligence required for dynamic, multi-objective missions in unstructured environments [1], [2], [3]. Achieving this shift demands UAVs that not only perceive and react but also reason, plan, and collaborate as intelligent, ecosystem-integrated agents.
Recent efforts to integrate Large Language Models (LLMs) into UAV autonomy highlight both the promise and the current limitations of this approach. Works such as REAL [4], UAV-VLN [5], and UAVs Meet LLMs [6] demonstrate how LLMs can parse natural language instructions or improve resilience through semantic reasoning. Aero-LLM [7] and UAV-CodeAgents [8] extend this by exploring distributed architectures and multi-agent planning. Similarly, Zhao and Lin [9] present general-purpose aerial agents through hardware-software co-design. However, across these approaches, the LLM typically functions as an isolated planner—a semantic parser or mission-level controller operating on curated state abstractions. These systems lack grounding in continuous sensor data, are not empowered to query external knowledge bases or APIs, and have limited capacity for peer-to-peer reasoning in multi-agent swarms. As a result, UAVs remain disconnected from the broader digital ecosystem and fall short of achieving true agentic intelligence [10], [11].
@@
We address this question by introducing the Agentic UAVs framework, a five-layer architecture (Perception, Reasoning, Action, Integration, and Learning). The key contributions are: (i) a structured integration of symbolic reasoning with continuous perception and control, (ii) an explicit Integration Layer for secure interaction with external APIs, databases, and peer agents using standardized protocols (MCP, ACP, A2A), and (iii) swarm cognition mechanisms enabling UAVs to negotiate tasks, share reasoning, and dynamically reallocate cognitive load. Experimental validation in simulated search-and-rescue scenarios shows sub-2-second decision latency, 92% contextual accuracy, and a 76% reduction in operator intervention compared to rule-based baselines.
@@ 2.1 From Programmed Automation to Reactive Control
Early UAV autonomy relied on precisely defined world models and deterministic execution, employing classical path-planning algorithms. This paradigm’s primary limitation was its brittleness, with performance degrading catastrophically when environments deviated from a priori models. Subsequent AI integration enhanced adaptivity. Vision-based methods, including Convolutional Neural Networks (CNNs) [2], [12], improved perception but lacked the semantic context for mission-level inference. Reinforcement Learning (RL) introduced a more significant advance, enabling UAVs to derive reactive control policies for tasks like obstacle avoidance [1]. However, RL’s optimization of low-dimensional state-action mappings, driven by scalar rewards, is fundamentally ill-equipped for long-horizon, multi-objective missions demanding abstract reasoning.
Similarly, swarm intelligence research has progressed from decentralized coordination to adaptive strategies, with works employing multi-agent reinforcement learning (MARL) for self-organizing swarm reconfiguration [3]. Despite this sophistication, these systems primarily optimize for coordinated motion and reactive redeployment, not for collaborative, high-level strategic reasoning or cognition.
@@ 2.2 The Emergence of Cognitive Reasoning with LLMs
The advent of LLMs has catalyzed a paradigm shift from purely reactive control to deliberative planning. Initial works demonstrate LLMs’ potential to parse high-level human commands into actionable sequences [9], [6]. For instance, Tagliabue et al. [4] show LLMs enhancing system resilience by interpreting diagnostic data to adapt control inputs, while UAV-VLN [5] focuses on grounding natural language instructions for navigation.
Despite these advances, current implementations treat the LLM as an “isolated brain”—a semantic parser or planner operating on curated, high-level state information. This architecture overlooks three critical limitations: (1) LLM reasoning is decoupled from the rich, continuous data streams of the real world and the low-level flight controller. (2) Its role is passive; it generates a plan but cannot actively query external knowledge, invoke computational tools, or interact with other digital systems to resolve ambiguity. (3) While distributed frameworks like Aero-LLM [7] address multi-agent LLM communication, they lack models for genuine collaborative problem-solving, such as distributed cognitive offloading or strategic negotiation. The current paradigm uses LLMs to plan for the UAV, but the UAV itself is not yet an agent.
@@ 3 Proposed Approach
Figure 1: Agentic UAVs: Five-Layer Architecture for General-Purpose Aerial Agents
@@ 3.2 The Five-Layer Architecture
  • 3. The Action Layer (Embodied and Ecosystem Execution): This layer operationalizes the Reasoning Layer’s plans across two domains: physical embodiment and digital ecosystem execution, making the UAV a networked digital actor.
@@
    • Ecosystem Execution (Tool-Actuation): Beyond flight, the Action Layer executes digital actions authorized by the reasoning engine. This grounds the agent in the broader digital world, allowing it to:
        Query External APIs: Access real-time data from weather services, aviation authorities (NOTAMs), or mapping platforms to enhance contextual awareness.
        Interact with Databases: Securely update mission-critical systems, such as logging an incident in a security database or updating medical triage records for first responders.
        Communicate and Report: Send structured alerts to emergency dispatch systems, messaging apps, or social media channels as part of an automated crisis response workflow.
        Execute Onboard Code: Run sandboxed code snippets (e.g., Python) for on-the-fly data analysis or custom image processing, extending its capabilities without pre-programming.
@@
    All actions, physical and digital, return structured feedback, ensuring the agent treats all interactions as contract-bound commitments subject to monitoring and reflection.
  • 4. The Integration Layer (Ecosystem and Swarm Execution): This layer extends the UAV’s agency beyond physical embodiment, acting as the gateway to digital ecosystems and multi-agent collaboration. It executes non-physical actions authorized by the Reasoning Layer and formalizes secure interaction with external services, databases, and peers.
    The Integration Layer provides twp key capabilities:
    • Tool-Actuation: Executes digital actions such as querying weather or NOTAM APIs, updating mission-critical databases (e.g., triage logs, security records), transmitting structured alerts to human operators or emergency systems, and running sandboxed code for on-the-fly data analysis.
@@
    • Protocol-Governed Communication: Ecosystem interaction is standardized through emerging multi-agent protocols:
@@
        MCP (Model Context Protocol): A framework by Anthropic that governs secure tool use by LLM agents, ensuring controlled context exchange and traceable reasoning.
@@
        ACP (Agent Communication Protocol): Structures UAV-to-cloud and UAV-to-operator messages, enabling interoperability and supervision.
@@
        Agent-to-Agent (A2A) Protocols: Facilitate peer-to-peer negotiation, task allocation, and consensus within UAV swarms for distributed cognition.
    By isolating ecosystem execution from flight control, this layer creates a clean boundary between physical action and digital cognition. This separation enhances modularity, security, and scalability across diverse UAV platforms.
    Unlike prior architectures where digital integration is ad hoc, this explicit layer makes ecosystem interaction a first-class capability. It transforms UAVs into not just mobile robots but active participants in socio-technical systems, capable of contributing knowledge, negotiating tasks, and augmenting human decision-making.
@@ 4.4.3 Decision-Making and Autonomy
(Cloud API) (network/queue) no HW req.
@@
(Local) (local I/O) (4B params) private
@@
Optimization insights: 70% speed improvement via local deployment; network infrastructure dominates cloud overhead; 4B-parameter model yields comparable reasoning at lower compute; hybrid Tier-1 (YOLO) + Tier-2 (Local LLM) enables near-real-time intelligent UAV operations.
@@ 5 Conclusion
This work proposed the Agentic UAVs framework, a five-layer architecture that integrates LLM-driven reasoning, perception, and ecosystem interaction to advance UAV autonomy beyond rule-based paradigms. Experimental validation in realistic SAR scenarios demonstrated significant gains in contextual understanding, decision-making, and autonomous intervention, confirming that computational overhead is offset by actionable intelligence and reduced operator dependence. While encouraging, challenges remain in minimizing latency, ensuring reliability under diverse mission conditions, and addressing safety and security in multi-agent operations. Future research should explore hybrid local–cloud deployments, scalable swarm cognition, and field validation to transition agentic UAVs from simulated prototypes to operationally robust aerial agents.

### v2

@@ 1 Introduction
Unmanned Aerial Vehicles (UAVs) are rapidly transitioning from remotely piloted platforms to autonomous agents capable of performing increasingly complex tasks in logistics, agriculture, defense, and disaster response. This transformation parallels advances in artificial intelligence (AI), where systems have evolved from rule-based automation to data-driven and adaptive reasoning. Despite progress in perception, planning, and control, UAV autonomy remains predominantly narrow—optimized for specific tasks but lacking the general-purpose intelligence required for dynamic, multi-objective missions in unstructured environments [zhang2023rl], [li2024cnn], [tian2025swarm]. Achieving this shift demands UAVs that not only perceive and react but also reason, plan, and collaborate as intelligent, ecosystem-integrated agents.
Recent efforts to integrate Large Language Models (LLMs) into UAV autonomy highlight both the promise and the current limitations of this approach. Works such as REAL [tagliabue2023real], UAV-VLN [saxena2025], and UAVs Meet LLMs [tian2025uavs] demonstrate how LLMs can parse natural language instructions or improve resilience through semantic reasoning. Aero-LLM [dharmalingam2025] and UAV-CodeAgents [sautenkov2025] extend this by exploring distributed architectures and multi-agent planning. Similarly, Zhao and Lin [zhao2025general] present general-purpose aerial agents through hardware-software co-design. However, across these approaches, the LLM typically functions as an isolated planner—a semantic parser or mission-level controller operating on curated state abstractions. These systems lack grounding in continuous sensor data, are not empowered to query external knowledge bases or APIs, and have limited capacity for peer-to-peer reasoning in multi-agent swarms. As a result, UAVs remain disconnected from the broader digital ecosystem and fall short of achieving true agentic intelligence [javaid2025], [innov82025].
@@
1.1 Contributions and Novelty
We address this question by introducing the Agentic UAVs framework, a novel five-layer architecture (Perception, Reasoning, Action, Integration, and Learning) that makes three key contributions:
  • Architectural Novelty: First five-layer architecture that tightly fuses LLM-driven symbolic reasoning with continuous sensor perception and low-level flight control, bridging the gap between deliberative planning and reactive execution.
  • Ecosystem Integration: Novel Integration Layer enabling UAVs to function as digital ecosystem participants through tool-calling, API interaction, and standardized multi-agent protocols (MCP, ACP, A2A)—transforming UAVs from isolated planners to networked cognitive agents.
  • Collaborative Swarm Cognition: First demonstration of distributed LLM-based reasoning for task negotiation, dynamic cognitive load allocation, and collaborative problem-solving among UAV swarms.
Experimental validation in simulated search-and-rescue scenarios demonstrates sub-2-second decision latency, 92% contextual accuracy, and 76% reduction in operator intervention compared to rule-based baselines.
@@ 2.1 From Programmed Automation to Reactive Control
Early UAV autonomy relied on precisely defined world models and deterministic execution, employing classical path-planning algorithms. This paradigm’s primary limitation was its brittleness, with performance degrading catastrophically when environments deviated from a priori models. Subsequent AI integration enhanced adaptivity. Vision-based methods, including Convolutional Neural Networks (CNNs) [li2024cnn], [redmon2016yolo], improved perception but lacked the semantic context for mission-level inference. Reinforcement Learning (RL) introduced a more significant advance, enabling UAVs to derive reactive control policies for tasks like obstacle avoidance [zhang2023rl]. However, RL’s optimization of low-dimensional state-action mappings, driven by scalar rewards, is fundamentally ill-equipped for long-horizon, multi-objective missions demanding abstract reasoning.
Similarly, swarm intelligence research has progressed from decentralized coordination to adaptive strategies, with works employing multi-agent reinforcement learning (MARL) for self-organizing swarm reconfiguration [tian2025swarm]. Despite this sophistication, these systems primarily optimize for coordinated motion and reactive redeployment, not for collaborative, high-level strategic reasoning or cognition.
@@ 2.2 The Emergence of Cognitive Reasoning with LLMs
The advent of LLMs has catalyzed a paradigm shift from purely reactive control to deliberative planning. Initial works demonstrate LLMs’ potential to parse high-level human commands into actionable sequences [zhao2025general], [tian2025uavs]. For instance, Tagliabue et al. [tagliabue2023real] show LLMs enhancing system resilience by interpreting diagnostic data to adapt control inputs, while UAV-VLN [saxena2025] focuses on grounding natural language instructions for navigation.
Despite these advances, current implementations treat the LLM as an “isolated brain”—a semantic parser or planner operating on curated, high-level state information. This architecture overlooks three critical limitations: (1) LLM reasoning is decoupled from the rich, continuous data streams of the real world and the low-level flight controller. (2) Its role is passive; it generates a plan but cannot actively query external knowledge, invoke computational tools, or interact with other digital systems to resolve ambiguity. (3) While distributed frameworks like Aero-LLM [dharmalingam2025] address multi-agent LLM communication, they lack models for genuine collaborative problem-solving, such as distributed cognitive offloading or strategic negotiation. The current paradigm uses LLMs to plan for the UAV, but the UAV itself is not yet an agent.
@@
Positioning Our Contribution. While recent works integrate LLMs for parsing commands or diagnostics, none achieve full agentic integration. Our framework is the first to combine (i) continuous perception-reasoning fusion, (ii) tool-calling for ecosystem interaction via standardized protocols (MCP, ACP, A2A), and (iii) collaborative swarm cognition within a unified five-layer architecture.
Table 1 positions our work against six representative recent approaches, highlighting what each lacks and how our framework addresses these gaps through architectural and functional novelty.
Table 1: Comparison with Recent LLM-UAV Approaches
Work Limitation Our Contribution
REAL [tagliabue2023real] LLM for diagnostics only; no tool use Tool-calling + API integration via Integration Layer
UAV-VLN [saxena2025] Vision-language navigation only General-purpose 5-layer architecture for diverse missions
UAVs+LLMs [tian2025uavs] Isolated LLM planner Perception-reasoning fusion with continuous feedback
Aero-LLM [dharmalingam2025] Multi-agent communication only Cognitive swarm via MCP & A2A protocols
Zhao [zhao2025general] No ecosystem integration Integration Layer with external APIs & databases
UAV-CodeAgents [sautenkov2025] Code generation focus Continuous perception loop + tool actuation
Summary: Our 5-layer architecture uniquely combines LLM tool-calling, ecosystem protocols, and swarm cognition in a cohesive framework, transforming UAVs from isolated planners to networked cognitive agents.
@@ 3 Proposed Approach
Figure 1: Agentic UAVs: Five-Layer Architecture for General-Purpose Aerial Agents. Layer 3 (Action) serves as the executor for physical and digital actions, while Layer 4 (Integration) provides the infrastructure for ecosystem management and multi-agent coordination.
@@ 3.2 The Five-Layer Architecture
  • 3. The Action Layer (Physical Embodiment and Tool Invocation): This layer operationalizes the Reasoning Layer’s plans by executing both physical and digital actions as atomic operations. It serves as the execution endpoint that bridges symbolic reasoning with concrete outcomes.
@@
    • Tool Invocation Interface: The Action Layer serves as the execution endpoint for tool calls authorized by the Reasoning Layer. When the reasoning engine decides to invoke an external capability (e.g., query weather API, log incident data), the Action Layer atomically executes this call via the Integration Layer’s infrastructure. This creates a clean separation of responsibilities: the Reasoning Layer decides what to do, the Action Layer executes the call, and the Integration Layer manages how tools are discovered, secured, and orchestrated.
      Example tool invocation flow:
      1. 1.
        Reasoning Layer: “I need weather data for coordinates (lat, lon)”
      2. 2.
        Action Layer: Invokes call_tool("weather.get_forecast", {lat, lon})
      3. 3.
        Integration Layer: Resolves service endpoint, handles authentication, executes via MCP protocol
      4. 4.
        Action Layer: Returns structured result to Reasoning Layer for next planning step
@@
    All actions, physical and digital, return structured feedback with execution status, ensuring the agent treats all interactions as contract-bound commitments subject to monitoring and reflection.
  • 4. The Integration Layer (Ecosystem Management and Multi-Agent Coordination): This layer provides the infrastructure that enables the UAV to interact with external digital services and collaborate with other agents. Unlike Layer 3, which executes individual tool calls as atomic actions, Layer 4 manages the lifecycle and orchestration of the UAV’s ecosystem interactions, making such operations possible, secure, and scalable.
    The Integration Layer provides five key capabilities:
    • Tool Registry & Service Discovery: Maintains a dynamic catalog of available tools (APIs, databases, services) with comprehensive metadata including schemas, endpoints, authentication requirements, rate limits, and availability status. Tools can be registered at runtime, enabling the UAV to adapt to new services without reprogramming. This registry acts as the single source of truth for all external capabilities accessible to the reasoning engine.
    • Execution Orchestration: Handles the mechanics of tool invocation beyond simple function calls, including:
        Authentication & authorization management (API keys, OAuth tokens, certificate validation)
        Request/response serialization, validation, and schema enforcement
        Retry logic, timeout handling, and circuit breaker patterns for resilience
        Result caching and memoization for efficiency
        Comprehensive telemetry, logging, and observability for debugging and audit trails
@@
    • Protocol-Governed Communication: Implements standardized multi-agent protocols to ensure interoperability and security:
@@
        MCP (Model Context Protocol): Governs secure tool use by LLM agents, ensuring controlled context exchange and traceable reasoning. Defines tool schemas, execution contracts, and safety boundaries for LLM-driven tool invocation.
@@
        ACP (Agent Communication Protocol): Structures UAV-to-cloud and UAV-to-operator messages for standardized telemetry, command, and supervision, enabling seamless integration with ground control systems.
@@
        Agent-to-Agent (A2A) Protocols: Enables peer-to-peer negotiation, task allocation, and consensus for distributed swarm intelligence (e.g., Contract Net Protocol, FIPA standards).
    • Security Mechanisms and Trust Model: The Integration Layer implements defense-in-depth security aligned with UAV threat models:
        Authentication & Authorization: MCP and ACP enforce mutual TLS (mTLS) with certificate-based authentication. API calls require signed JWTs with scoped permissions, preventing unauthorized tool invocation and command injection. OAuth 2.0 integration enables fine-grained access control for external services.
        Encrypted Communication: All inter-agent (A2A) and UAV-to-cloud (ACP) traffic uses AES-256-GCM encryption over TLS 1.3. Message integrity is ensured via HMAC-SHA256 signatures, detecting tampering and replay attacks.
        LLM-Specific Safeguards: MCP implements prompt injection detection, output sanitization, and tool schema validation to mitigate adversarial inputs. Execution sandboxing limits LLM agent capabilities to predefined tool boundaries.
        Threat Mitigation: The design addresses GPS spoofing (cross-validation with inertial sensors), denial-of-service (rate limiting, circuit breakers), and man-in-the-middle attacks (certificate pinning, mutual authentication).
    • Multi-Agent Coordination: Facilitates swarm behaviors and collaborative decision-making:
        Task decomposition and allocation among multiple UAVs based on capability and availability
        Spatial deconfliction and conflict resolution to prevent collisions and resource contention
        Shared situational awareness through distributed scene graphs and knowledge synchronization
        Consensus protocols for joint decision-making in contested or uncertain environments
    This layer transforms the UAV from an isolated autonomous system into a networked digital actor. By separating ecosystem management (Layer 4) from action execution (Layer 3), the architecture achieves modularity, enables dynamic capability extension, and supports both single-UAV and swarm deployments with minimal reconfiguration. Unlike prior architectures where digital integration is ad hoc, this explicit layer makes ecosystem interaction a first-class architectural capability, positioning UAVs as active participants in socio-technical systems.
@@ 4.4.3 Decision-Making and Autonomy
(Cloud API) (network/queue) (ARR: 92%) no HW req.
@@
(Local) (local I/O) (ARR: 79%) private
@@
Optimization insights: 70% speed improvement via local deployment; network infrastructure dominates cloud overhead; 4B-parameter model offers operationally acceptable performance (79% ARR vs. 92% for GPT-4) with favorable speed-accuracy trade-off at significantly lower computational cost; hybrid Tier-1 (YOLO) + Tier-2 (Local LLM) enables near-real-time intelligent UAV operations.
@@ 5 Conclusion
This work proposed the Agentic UAVs framework, a five-layer architecture that integrates LLM-driven reasoning, perception, and ecosystem interaction to advance UAV autonomy beyond rule-based paradigms. Experimental validation in realistic SAR scenarios demonstrated significant gains in contextual understanding, decision-making, and autonomous intervention, confirming that computational overhead is offset by actionable intelligence and reduced operator dependence.
5.1 Limitations and Future Work
Simulation-Based Validation. The current evaluation relies on high-fidelity Gazebo-based simulation rather than real-world deployment. This choice was driven by safety considerations (avoiding risks during agentic UAV development), cost efficiency, and experimental reproducibility. While simulation provides controlled conditions for systematic performance analysis, it introduces a
textitsim-to-real gap in dynamics modeling, sensor noise, and environmental variability. Real-world factors such as wind turbulence, GPS degradation, and communication latency may affect system performance differently than simulated counterparts.
Future Directions. Immediate priorities include field trials in controlled outdoor environments to validate the framework’s robustness under real conditions, hybrid local–cloud deployments to optimize latency-accuracy trade-offs, and scalable swarm cognition for multi-UAV collaboration. Longer-term research should address safety assurance for LLM-driven decision-making, regulatory compliance for autonomous UAV operations, and integration with standardized UTM (Unmanned Traffic Management) systems. These steps are essential to transition agentic UAVs from simulated prototypes to operationally robust aerial agents.
Author Contribution Statement
The intellectual property and conceptual foundation of this work, including the definition of the Agentic UAVs and Flying Agents frameworks, originate from Prof. Anis Koubaa, who also served as the main author of the manuscript. Khaled Gabr contributed to the development and implementation of the ROS 2-based validation framework, as well as the design and execution of simulation experiments and the preparation of the results section. Both authors reviewed and approved the final version of the paper.